OpenStudy (cutiecomittee123):

Solve on an interval of [0,2pi) 2secx+5=1 I solved it down to sec x = -1/2 I know that sec x = 1/cos x but now I am stuck please help!

OpenStudy (cutiecomittee123):

@zepdrix can you help?

OpenStudy (baru):

2secx=-4 secx=-4/2 secx=-2

OpenStudy (pawanyadav):

Sec x=-2. So cos x=-1/2

OpenStudy (cutiecomittee123):

oh yeah i did the calculations wrong. I got it thanks!

OpenStudy (cutiecomittee123):

so its 2pi/3 and 4pi/3

OpenStudy (pawanyadav):

Right